### v2

# CLASP: An online learning algorithm for Convex Losses And Squared Penalties

Abstract
We study Constrained Online Convex Optimization (COCO), where a learner chooses actions iteratively, observes both unanticipated convex loss and convex constraint, and accumulates loss while incurring penalties for constraint violations. We introduce CLASP (Convex Losses And Squared Penalties), an algorithm that minimizes cumulative loss together with squared constraint violations. Our analysis departs from prior work by fully leveraging the firm non-expansiveness of convex projectors, a proof strategy not previously applied in this setting. For convex losses, CLASP achieves regret $O\left(T^{\max\{\beta,1-\beta\}}\right)$ and cumulative squared penalty $O\left(T^{1-\beta}\right)$ for any $\beta\in(0,1)$. Most importantly, for strongly convex problems, CLASP provides the first logarithmic guarantees on both regret and cumulative squared penalty. In the strongly convex case, the regret is upper bounded by $O(\log T)$ and the cumulative squared penalty is also upper bounded by $O(\log T)$.
1 Introduction
We consider a setting where at each iteration $t\in\{1,2,3,\dots\}$, a learner selects an action $x_{t}$ from a convex set $\mathcal{K}\subset\mathbb{R}^{n}$, then a loss function $f_{t}$ is revealed, and the learner incurs loss $f_{t}(x_{t})$. The learner’s goal is to perform nearly as well dynamically as the best fixed action in hindsight. i.e, keep the regret
$\text{Regret}_{T}=\sum_{t=1}^{T}f_{t}(x_{t})-\sum_{t=1}^{T}f_{t}(x^{\star}_{T}).$ (1)
growing sublinearly in $T$, so that, asymptotically, the average loss of the learner is no worse than the optimal fixed action, $x_{T}^{\star}=\underset{x\in\mathcal{K}}{\arg\min}\sum_{t=1}^{T}f_{t}(x)$.
Online Convex Optimization (OCO) is the special case of Online Learning in which the action set $\mathcal{K}$ is convex, and the loss functions $f_{t}$ are convex. The OCO setup has been extensively studied over the years (Shalev-Shwartz and others, 2012; Hazan and others, 2016; Zinkevich, 2003; Duchi et al., 2011; Bubeck et al., 2012; Hazan and Kale, 2012; Hazan and Minasyan, 2020; Hazan and Singh, 2021). OCO assumes that operational constraints on the possible actions are static in time, being fully captured by the fixed set $\mathcal{K}$.
In many applications, however, operational constraints do change at each iteration: at time $t$, the action $x_{t}$ must satisfy not only $x_{t}\in\mathcal{K}$ but also an extra constraint $g_{t}(x_{t})\leq 0$, with $g_{t}$ a convex function. The challenge here is that the learner must choose an action $x_{t}$ also before knowing the constraint function $g_{t}$. This more difficult setup generalizes OCO and is known as Constrained Online Convex Optimization (COCO).
To conclude, the COCO learner wishes not only to bound the regret (1), now with
$x_{T}^{\star}=\underset{x\in\mathcal{K}\cap{\mathcal{C}}_{T}}{\arg\min}\sum_{t=1}^{T}f_{t}(x),\quad{\mathcal{C}}_{T}=\bigcap_{t=1}^{T}C_{t},\quad C_{t}=\{x\,\colon\,g_{t}(x)\leq 0\},$ (2)
but also wishes to bound some notion of cumulative constraint violation (CCV), as detailed in the next section.
1.1 Related Work
Early work on COCO (Mahdavi et al., 2012; Jenatton et al., 2016; Yuan and Lamperski, 2018; Yu and Neely, 2020) focused on soft constraint violations, allowing the learner to compensate over time by balancing positive and negative violations. In those settings the performance is measured by the cumulative constraint violation $\text{CCV}_{T}=\sum_{t=1}^{T}g_{t}(x_{t})$. However, this metric has a critical weakness: The sum may become negative even if many rounds incur positive violations, thereby masking the severity or frequency of actual constraint breaches.
Such arithmetic compensation is often unsuitable in practice, as many applications demand a more direct and monotonic measure of constraint violation. We thus focus on the hard metrics
$\text{CCV}_{T,1}=\sum_{t=1}^{T}g_{t}^{+}(x_{t}),\quad\text{or}\quad\text{CCV}_{T,2}=\sum_{t=1}^{T}\left(g_{t}^{+}(x_{t})\right)^{2},$ (3)
where $g_{t}^{+}(x)=\max\{0,g_{t}(x)\}$. Under these definitions, violations cannot be offset over time: these cumulative sums are nondecreasing in $T$. Both metrics are of practical relevance, and authors generally commit to one as their principal violation measure. A natural question is whether bounds on $\text{CCV}_{T,1}$ can imply (in some way) bounds on $\text{CCV}_{T,2}$. Indeed, one might attempt to reduce to a $\text{CCV}_{T,2}$-style problem by redefining the constraint functions as $\tilde{g}_{t}=(g_{t}^{+})^{2}$. However, that change can invalidate critical assumptions (e.g., Slater’s condition, convexity, or Lipschitz bounds) originally required on each $g_{t}$. Therefore, in this paper we directly target $\text{CCV}_{T,2}$ to derive explicit upper bounds for it.
We now review the closest work in COCO, organized by whether the constraints are static ($g_{t}=g$ for all $t$) or dynamic ($g_{t}$ may vary adversarially). Our focus is on the dynamic case, but it is useful to contrast with the static regime first.
Static constraints.
As the focus of this paper is on dynamic constraints, we restrict ourselves here for brevity to convex loss functions, omitting the results for strongly convex ones. The early work of Mahdavi et al. (2012) uses a regularized Lagrangian update and achieves $\mathrm{Regret}_{T}=O(\sqrt{T})$ and soft violation $\mathrm{CCV}_{T}=O(T^{3/4})$Jenatton et al. (2016) refine this approach via adaptive weightings of primal and dual step sizes, obtaining $\mathrm{Regret}_{T}=O\!\bigl(T^{\max\{\beta,1-\beta\}}\bigr)$ and $\mathrm{CCV}_{T}=O\!\bigl(T^{1-\beta/2}\bigr)$ for any $\beta\in(0,1)$. Yu and Neely (2020) push further by bounding $\mathrm{CCV}_{T}=O(1)$ (constant violation) while preserving $O(\sqrt{T})$ regret. Yi et al. (2021), in turn, address hard violations by showing $\mathrm{Regret}_{T}=O\!\bigl(T^{\max\{\beta,1-\beta\}}\bigr)$ and $\mathrm{CCV}_{T,1}=O\bigl(T^{(1-\beta)/2}\bigr)$. Finally, Yuan and Lamperski (2018) focus directly on the squared‐penalty metric $\mathrm{CCV}_{T,2}$, proving $\mathrm{Regret}_{T}=O\!\bigl(T^{\max\{\beta,1-\beta\}}\bigr)$ and $\mathrm{CCV}_{T,2}=O(T^{1-\beta})$. Their work penalizes large violations more heavily, preventing cancellation that occurs under soft violation metrics.
Dynamic constraints.
When the constraints $g_{t}$ may vary arbitrarily with $t$, the problem becomes significantly more challenging.
Guo et al. (2022) introduce the Rectified Online Optimization (RECOO) algorithm, which exploits a first‐order approximation of the regularized Lagrangian at each round. For convex losses, RECOO achieves $\text{Regret}_{T}\leq O(\sqrt{T})$ and $\text{CCV}_{T,1}\leq O\left(T^{\frac{3}{4}}\right)$. In the strongly convex case, it improves to $\text{Regret}_{T}\leq O(\log T)$ and $\text{CCV}_{T,1}\leq O(\sqrt{T\log T})$.
Yi et al. (2023) extend these ideas in a distributed setting, with results that can be transposed to the centralized regime. For convex losses, they establish $\text{Regret}_{T}\leq O\left(T^{\max\{\beta,1-\beta\}}\right)$ and $\text{CCV}_{T,1}\leq\left(T^{1-\beta/2}\right)$ for any $\beta\in(0,1)$. For strongly convex losses, their guarantees become $\text{Regret}_{T}\leq O\left(T^{\beta}\right)$ and $\text{CCV}_{T,1}\leq\left(T^{1-\beta/2}\right)$.
A major step forward is due to Sinha and Vaze (2024), who introduced the Regret Decomposition Inequality, an elegant analytical tool that extends the drift-plus-penalty framework (Neely, 2010). They show that suitably modified AdaGrad variants can achieve $\text{Regret}_{T}\leq O\left(\sqrt{T}\right)$ and, for the first time, $\text{CCV}_{T,1}\leq O\left(\sqrt{T}\log T\right)$ for convex loss functions. For strongly convex functions, the results are $\text{Regret}_{T}\leq O\left(\log T\right)$ and $\text{CCV}_{T,1}\leq O\left(\sqrt{T\log T}\right)$, with the CCV bound improved to $\text{CCV}_{T,1}\leq O(\log T)$ for the specialized setting of non-negative regrets.
Building on this line Vaze and Sinha (2025), propose the Switch algorithm, which explicitly leverages the geometry of the constraint sets $C_{t}$ (2). Although the strongly convex case is not explicitly addressed, for convex loss functions Switch guarantees $\text{Regret}_{T}\leq O(\sqrt{T})$ and $\text{CCV}_{T,1}\leq O(\sqrt{T}\log T)$ across all instances, the novelty being that the bound on CCV can drop to $O(1)$ for special geometrical instances of $C_{t}$. Our proposed algorithm, CLASP, bears certain similarities with Algorithm 2 in (Vaze and Sinha, 2025), as we elaborate in the next section.
To conclude, we highlight an interesting line of COCO research on low-complexity, projection-free algorithms, which avoid projection onto $\mathcal{K}$. Representative work is (Garber and Kretzu, 2024; Sarkar et al., 2025; Wang et al., 2025; Lu et al., 2025). For example, the recent work Lu et al. (2025) exploits a separation oracle to achieve $\text{Regret}_{T}\leq O\left(\sqrt{T}\right)$ and $\text{CCV}_{T,1}\leq O\left(\sqrt{T}\log T\right)$ for convex loss functions; and $\text{Regret}_{T}\leq O\left(\log T\right)$ and $\text{CCV}_{T,1}\leq O\left(\sqrt{T\log T}\right)$ for strongly convex ones.
1.2 Contributions
We propose CLASP, an online COCO algorithm for dynamic constraints with Convex Losses And Squared Penalties, where the violation metric of interest is $\mathrm{CCV}_{T,2}$ (3). Our contributions can be summarized as follows:
  • Strongly convex losses. CLASP achieves $\mathrm{Regret}_{T}\leq O(\log T)$ and $\mathrm{CCV}_{T,2}\leq O(\log T)$. To the best of our knowledge, this is the first result to guarantee logarithmic bounds on both regret and squared violations. The closest prior work, (Sinha and Vaze, 2024), establishes $\mathrm{Regret}_{T}\leq O(\log T)$ and $\mathrm{CCV}_{T,1}\leq O(\log T)$, but only in the restricted setting of non-negative regrets. CLASP removes this limitation, providing guarantees without assumptions on the sign of the regret.
  • Convex losses. For general convex losses, CLASP guarantees
    $\mathrm{Regret}_{T}\leq O\!\bigl(T^{\max\{\beta,1-\beta\}}\bigr),\quad\mathrm{CCV}_{T,2}\leq O(T^{1-\beta}),\qquad\forall\,\beta\in(0,1).$
    This matches the rates of (Yuan and Lamperski, 2018), whose results were derived for static constraints; CLASP extends them to the dynamic regime.
  • Algorithmic simplicity. Each CLASP iteration consists of a gradient step with respect to the latest loss, followed by a single projection onto the current feasible set ${\mathcal{K}}\cap C_{t}$. In contrast, Algorithm 2 of (Vaze and Sinha, 2025) requires two projections per iteration, one onto the full historical intersection ${\mathcal{K}}\cap\bigcap_{\tau=1}^{t}C_{\tau}$. Thus CLASP is more memory-efficient, avoiding the need to retain past constraints. Moreover, the analyses of the two methods differ fundamentally, as discussed in Sections 5 and 6.
  • Analytical novelty. Our analysis relies on the firm non-expansiveness (FNE) of projections onto closed convex sets—a stronger property than the standard non-expansiveness typically used. Exploiting FNE allows a clean modular proof structure, where regret and constraint violation are analyzed separately. This modularity also makes the analysis more extensible, e.g., to multiple dynamic constraints or persistent constraints (Section 6).
Relevance to Machine Learning Practice.
Recent empirical and theoretical work suggests that strong convexity and sharper penalties for violations are practically meaningful in constrained ML settings. For example, Wang et al. (2025) shows that under strong convexity of regret, both regret and constraint violations drop substantially, and experiments in that paper confirm that algorithms exploiting strong convexity perform better on real datasets. The authors of Ma et al. (2025) empirically validate that in resource allocation problems with hard constraints plus regularization, faster rates (logarithmic regret) are achievable. The work of Banerjee et al. (2023), finds that many practical deep models satisfy restricted strong convexity, and fast (geometric) convergence is observed in training. These works show that assumptions like strong convexity are not merely theoretical and that penalties, and constraints (Ramirez et al. (2025)), matter in real ML systems. However, prior work usually measures constraint violation linearly (or counts violations) rather than providing log‐rate guarantees on both regret and the severity of violations (e.g., squared penalties).
2 Preliminaries
We summarize the technical tools used in the analysis of CLASP and state our assumptions.
Projection operators and distance functions.
For a non-empty, closed convex set $S\subset\mathbb{R}^{n}$, we let $P_{S}\,\colon\,\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be the associated orthogonal projection operator. Thus, $P_{S}(u)$ is the projection of the point $u$ onto the set $S$ and denotes the point in $S$ that is closest to $u$ with respect to the Euclidean norm $\left\|\cdot\right\|$. Such projectors are firmly non-expansive (FNE) operators, which means they satisfy the inequality
$\left\|P_{S}(u)-P_{S}(v)\right\|^{2}\leq\left\|u-v\right\|^{2}-\left\|(u-P_{S}(u))-(v-P_{S}(v))\right\|^{2},$ (4)
for all $u,v\in\mathbb{R}^{n}$ (see, e.g., Bauschke and Combettes (2017, Proposition 4.8)). Property (4) implies the popular non-expansiveness (NE) property, $\left\|P_{S}(u)-P_{S}(v)\right\|\leq\left\|u-v\right\|$.
We let $d_{S}\,\colon\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ denoted the associated distance function, $d_{S}(u)=\min\left\{\left\|v-u\right\|\,\colon\,v\in S\right\}$. Thus, $d_{S}(u)=\left\|u-P_{S}(u)\right\|$ and, from (4),
$\left\|P_{S}(u)-v\right\|^{2}\leq\left\|u-v\right\|^{2}-d_{S}(u)^{2},$ (5)
for all $u\in\mathbb{R}^{n}$ and $v\in S$.
Finally, the function $d_{S}$ is Lipschitz continuous with constant 1, that is, $\left|d_{S}(u)-d_{S}(v)\right|\leq\|u-v\|$ for all $u,v\in\mathbb{R}^{n}$ (see, e.g., Bauschke and Combettes (2017, page 59, Chapter 4)).
Convex and strongly convex functions.
Let $h\,\colon\,\mathbb{R}^{n}\rightarrow\,\mathbb{R}\cup\{+\infty\}$ be an extended real-valued function. It is said to be proper if its domain, $\text{dom }h=\{x\in\mathbb{R}^{n}\,\colon\,h(x)<+\infty\}$, is a non-empty set; it is closed if each lower level set $\{x\in\mathbb{R}^{n}\,\colon\,h(x)\leq\alpha\}$ at height $\alpha\in\mathbb{R}$ is a closed set; and it is convex if there exists an $m\geq 0$ such that
$h\left((1-\lambda)u+\lambda v\right)\leq(1-\lambda)h(u)+\lambda h(v)-\frac{m}{2}\lambda(1-\lambda)\left\|u-v\right\|^{2},$ (6)
for all $\lambda\in(0,1)$ and $u,v\in\mathbb{R}^{n}$. If $m$ can be taken to be strictly positive, then $h$ is said to be $m$-strongly convex.
If $h$ is a proper, closed, and convex function, then
$h(v)\geq h(u)+\langle\nabla h(u),v-u\rangle+\frac{m}{2}\left\|v-u\right\|^{2},$ (7)
for all $u,v\in\mathbb{R}^{n}$, where, from now on, $\nabla h(u)$ denotes a sub-gradient of $h$ at $u$. ($\langle\cdot,\cdot\rangle$ is the usual inner-product). If, furthermore, $h$ is $m$-strongly convex, then it has a unique global minimizer, say, $u^{\star}$, (see, e.g., Bauschke and Combettes (2017, Corollary 11.16)); the zero vector is then a sub-gradient of $h$ at $u^{\star}$, implying via (7) that
$h(v)\geq h(u^{\star})+\frac{m}{2}\left\|v-u^{\star}\right\|^{2}$ (8)
for all $v\in\mathbb{R}^{n}$.
Assumptions.
Throughout, we impose the following standard COCO assumptions:
Assumption 1.
The action set $\mathcal{K}$ is a non-empty, compact convex subset of $\mathbb{R}^{n}$. It follows that the diameter of $\mathcal{K}$ is upper-bounded by some constant, say, $D\geq 0$, which entails $\left\|u-v\right\|\leq D$ for all $u,v\in{\mathcal{K}}$.
Assumption 2.
The loss functions $f_{t}:\mathbb{R}^{n}\to\mathbb{R}$ are convex for all $t\geq 1$, with the magnitude of their sub-gradients bounded by $L$, when evaluated at points in $\mathcal{K}$: $\left\|\nabla f_{t}(u)\right\|\leq L$ for all $u\in{\mathcal{K}}$ and $t\geq 1$.
Assumption 3.
The constraint functions $g_{t}:\mathbb{R}^{n}\to\mathbb{R}$ are convex for all $t\geq 1$, with the magnitude of their sub-gradients bounded by $L$, when evaluated at points in $\mathcal{K}$: $\left\|\nabla g_{t}(u)\right\|\leq L$ for all $u\in{\mathcal{K}}$ and $t\geq 1$.
Moreover, ${\mathcal{K}}\cap\bigcap_{t=1}^{T}C_{t}$ is non-empty for all $T\geq 1$ (where $C_{t}=\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\}$), so as to ensure the existence of the regret comparator $x_{T}^{\star}$—recall (2))
Assumptions 2 and 3 imply that
$|f_{t}(v)-f_{t}(u)|\leq L\left\|v-u\right\|\quad\text{and}\quad|g_{t}(v)-g_{t}(u)|\leq\left\|v-u\right\|$ (9)
for all $u,v\in\mathcal{K}$ and $t\geq 1$ (as a consequence, e.g., of the mean-value theorem (Hiriart-Urruty and Lemaréchal, 1993, Theorem 2.3.3)).
Assumption 4.
The sequence (of step-sizes) $(\eta_{t})_{t\geq 1}$ satisfies $0<\eta_{t+1}\leq\eta_{t}$ for all $t\geq 1$. Hence, there exists $\theta>0$ such that $\eta_{t}^{2}\leq\theta\eta_{t}$ for all $t\geq 1$ (e.g., take $\theta=\eta_{1}$).
3 The CLASP algorithm
CLASP is a conceptually simple algorithm that, at each iteration (after the first), takes a gradient step with respect to the most recently observed loss function, and then projects onto the most recently observed constraint function, see Algorithm 1.
Algorithm 1 CLASP
action set $\mathcal{K}$, horizon $T\geq 1$, step-sizes $\eta_{t}$ for $1\leq t\leq T-1$
Choose $x_{1}\in\mathcal{K}$ and observe $f_{1},g_{1}$
Accumulate loss $f_{1}(x_{1})$ and penalty $(g_{1}^{+}(x_{1}))^{2}$
for $t=1,\dots,T-1$ do
  Choose $x_{t+1}=\mathcal{P}_{{\mathcal{K}}_{t}}\left(x_{t}-\eta_{t}\nabla f_{t}(x_{t})\right)$ and observe $f_{t+1},g_{t+1}$
  Accumulate loss $f_{t+1}(x_{t+1})$ and penalty $(g_{t+1}^{+}(x_{t+1}))^{2}$
end for
Here, $\nabla f_{t}(x_{t})$ represents a sub-gradient of $f_{t}$ at $x_{t}$. The projection step in CLASP is onto the set ${\mathcal{K}}_{t}={\mathcal{K}}\cap C_{t}$, with $C_{t}=\left\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\right\}$ (as defined in (2)). Note that the set ${\mathcal{K}}_{t}$ is non-empty, as per Assumption 3, and that this projection, typically realized as a convex optimization problem, sets CLASP apart from the projection-free methods in Section $1.$ The step-sizes $\eta_{t}$ are chosen according to whether the loss functions are convex or strongly convex and are specified further ahead in Section 5. We write $\widehat{x}_{t+1}=x_{t}-\eta_{t}\nabla f_{t}(x_{t})$, and rewrite CLASP as $x_{t+1}=p_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})$ for $t\geq 1$.
We now analyze CLASP, finding upper-bounds for both $\text{Regret}_{T}$ (see (1) and (2)) and the hard constraint violation $\text{CCV}_{T,2}$ (see (3)).
4 Bounding $\text{CCV}_{T,2}$
We start by finding a generic bound on $\text{CCV}_{T,2}$ that is phrased in terms of the length of the step-sizes $\sum_{t=1}^{T}\eta_{t}$. Our bound, the end result of a progression of three lemmas, is stated precisely in Lemma 3. The lemmas hold under Assumptions 1 to  4, and their proofs are provided in the Appendix.
Lemma 1.
There holds $\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}\leq O\left(\sum_{t=1}^{T}\eta_{t}\right)$.
Lemma 2.
There holds $\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(x_{t})^{2}\leq O\left(\sum_{t=1}^{T}\eta_{t}\right)$.
Lemma 3.
There holds $\text{CCV}_{T,2}\leq O\left(\sum_{t=1}^{T}\eta_{t}\right)$.
5 Analysis of CLASP
Having assembled the necessary elements, we are ready to fully analyze CLASP. The analysis is reported in Theorem 1 for convex losses, and in Theorem 2 for strongly convex ones.
Convex losses.
We suppose that assumptions 1 to 4 hold. Let $\beta$ be any desired value in $(0,1)$ and, accordingly, set the step-sizes to
$\eta_{t}=1/t^{\beta},\quad\text{for }t\geq 1,$ (10)
which complies with Assumption 4. The following Theorem 1 bounds the regret and the CCV of CLASP for this setting.
Theorem 1 (Convex losses).
CLASP achieves
$\text{Regret}_{T}\leq O(T^{\max\{\beta,1-\beta\}})\quad\text{and}\quad\text{CCV}_{T,2}\leq O\left(T^{1-\beta}\right).$
Proof.
We start by analyzing the $\text{Regret}_{T}$. Express the projection step of CLASP as
$x_{t+1}=\underset{u\in\mathbb{R}^{n}}{\arg\min}\ h_{t}(u),$
where
$h_{t}(u)=\langle\nabla f_{t}(x_{t}),u-x_{t}\rangle+\frac{1}{2\eta_{t}}\|u-x_{t}\|^{2}+\delta_{{\mathcal{K}}_{t}}(u),$
with $\delta_{{\mathcal{K}}_{t}}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ denoting the indicator function of the convex set ${\mathcal{K}}_{t}$ (i.e., $\delta_{{\mathcal{K}}_{t}}(u)=0$ if $u\in{\mathcal{K}}_{t}$, and $\delta_{{\mathcal{K}}_{t}}(u)=+\infty$, if $u\notin{\mathcal{K}}_{t}$).
Note that $h_{t}$ is a proper, closed, and $\sigma$-strongly convex function with modulus $\sigma=1/\eta_{t}$ and global minimizer $x_{t+1}$. Thus, from (8), we have
$h_{t}(x_{T}^{\star})\geq h_{t}(x_{t+1})+\frac{1}{2\eta_{t}}\left\|x_{T}^{\star}-x_{t+1}\right\|^{2}.$ (11)
Observing that both $x_{t+1}$ and $x_{T}^{\star}$ are points in ${\mathcal{K}}_{t}$ (which implies $\delta_{{\mathcal{K}}_{t}}(x_{t+1})=\delta_{{\mathcal{K}}_{t}}(x_{T}^{\star})=0$), we can re-arrange (11) as
$\frac{1}{2\eta_{t}}\|x_{t+1}-x_{T}^{\star}\|^{2}$ $\leq\frac{1}{2\eta_{t}}\|x_{t}-x_{T}^{\star}\|^{2}+\langle\nabla f_{t}(x_{t}),x_{T}^{\star}-x_{t}\rangle-\frac{1}{2\eta_{t}}\|x_{t}-x_{t+1}\|^{2}$ (12)
$+\langle\nabla f_{t}(x_{t}),x_{t}-x_{t+1}\rangle.$
We now use the easily-checked fact
$\max\left\{-\frac{1}{2\eta}\left\|u\right\|^{2}+\langle v,u\rangle\,\colon\,u\in\mathbb{R}^{n}\right\}=\frac{\eta}{2}\left\|v\right\|^{2},$
which holds for all $\eta>0$ and $v\in\mathbb{R}^{n}$, and amounts simply to compute the peak value of the concave quadratic function inside the braces. Using this fact with $\eta=\eta_{t}$ and $v=\nabla f_{t}(x_{t})$, we can bound the sum of the last two terms on the right-hand side of (12), thereby arriving at
$\frac{1}{2\eta_{t}}\|x_{t+1}-x_{T}^{\star}\|^{2}\leq\frac{1}{2\eta_{t}}\|x_{t}-x_{T}^{\star}\|^{2}+\langle\nabla f_{t}(x_{t}),x_{T}^{\star}-x_{t}\rangle+\frac{\eta_{t}}{2}\left\|\nabla f_{t}(x_{t})\right\|^{2}.$ (13)
The convexity of $f_{t}$ implies $\langle\nabla f_{t}(x_{t}),x_{T}^{\star}-x_{t}\rangle\leq f(x_{T}^{\star})-f(x_{t})$ (e.g., see (7) with $m=0$), which, when plugged in (13), yields
$f_{t}(x_{t})-f_{t}(x_{T}^{\star})\leq\frac{1}{2\eta_{t}}\|x_{t}-x_{T}^{\star}\|^{2}-\frac{1}{2\eta_{t}}\|x_{t+1}-x_{T}^{\star}\|^{2}+\frac{\eta_{t}}{2}\|\nabla f_{t}(x_{t})\|^{2}.$ (14)
For $t=1$, inequality (14) implies
$f_{1}(x_{1})-f_{1}(x_{T}^{\star})\leq\frac{1}{2\eta_{1}}D^{2}-\frac{1}{2\eta_{1}}\left\|x_{2}-x_{T}^{\star}\right\|^{2}+\frac{\eta_{1}}{2}\left\|\nabla f_{1}(x_{1})\right\|^{2},$ (15)
where Assumption 1 was used to bound the first term in the right-hand side of (14). For $t\geq 2$, we have
$\frac{1}{2\eta_{t}}\|x_{t}-x_{T}^{\star}\|^{2}$ $=\frac{1}{2\eta_{t-1}}\|x_{t}-x_{T}^{\star}\|^{2}+\left(\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\right)\|x_{t}-x_{T}^{\star}\|^{2}$
$\leq\frac{1}{2\eta_{t-1}}\|x_{t}-x_{T}^{\star}\|^{2}+\left(\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\right)D^{2},$ (16)
where the last inequality uses Assumptions 1 and 4.
Plugging (16) in (14) gives
$f_{t}(x_{t})-f_{t}(x_{T}^{\star})$ $\leq\frac{1}{2\eta_{t-1}}\|x_{t}-x_{T}^{\star}\|^{2}-\frac{1}{2\eta_{t}}\|x_{t+1}-x_{T}^{\star}\|^{2}+\frac{\eta_{t}}{2}\|\nabla f_{t}(x_{t})\|^{2}$ (17)
$+\left(\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\right)D^{2},$
which is valid for $t\geq 2$. Finally, summing (15) to (17) (instantiated from $t=2$ to $T$) yields
$\text{Regret}_{T}$ $\leq\frac{1}{2\eta_{T}}D^{2}+\sum_{t=1}^{T}\frac{\eta_{t}}{2}\left\|\nabla f_{t}(x_{t})\right\|^{2}$
$\leq\frac{T^{\beta}D^{2}}{2}+\frac{L^{2}}{2}\sum_{t=1}^{T}\frac{1}{t^{\beta}}$
$\leq O(T^{\max\{\beta,1-\beta\}}),$
where the second inequality follows from Assumption 2 and (10).
We now turn to the analysis of $\text{CCV}_{2,T}$: combining Lemma 3 with (10) gives directly $\text{CCV}_{T,2}\leq O(T^{1-\beta})$ and concludes the proof. ∎
Strongly convex losses.
We suppose assumptions 1 to 4 are still in force and further assume that there exists $m>0$ such that each convex loss $f_{t}$ is $m$-strongly convex for $t\geq 1$. The step-sizes are set to
$\eta_{t}=1/(mt),\quad\text{for }t\geq 1,$ (18)
thereby satisfying Assumption 4.
Theorem 2 (Strongly convex losses).
CLASP achieves
$\text{Regret}_{T}\leq O(\log T)\quad\text{and}\quad\text{CCV}_{T,2}\leq O(\log T).$
Proof.
Looking first at $\text{Regret}_{T}$ and replaying the opening moves in the proof of Theorem 1, we may assume that inequality (13) holds, an inequality that we reproduce here for convenience of the reader:
$\frac{1}{2\eta_{t}}\|x_{t+1}-x_{T}^{\star}\|^{2}\leq\frac{1}{2\eta_{t}}\|x_{t}-x_{T}^{\star}\|^{2}+\langle\nabla f_{t}(x_{t}),x_{T}^{\star}-x_{t}\rangle+\frac{\eta_{t}}{2}\left\|\nabla f_{t}(x_{t})\right\|^{2}.$ (19)
The strong convexity of $f_{t}$ implies, via (7),
$\langle f_{t}(x_{t}),x_{T}^{\star}-x_{t}\rangle\leq f_{t}(x_{T}^{\star})-f_{t}(x_{t})-(m/2)\left\|x_{t}-x_{T}^{\star}\right\|^{2},$
which, when inserted in (19), gives
$f_{t}(x_{t})-f_{t}(x_{T}^{\star})\leq\frac{m(t-1)}{2}\left\|x_{t}-x_{T}^{\star}\right\|^{2}-\frac{mt}{2}\left\|x_{t+1}-x_{T}^{\star}\right\|^{2}+\frac{\eta_{t}}{2}\left\|\nabla f_{t}(x_{t})\right\|^{2}$ (20)
(where we used (18) to replace $\eta_{t}$ in some terms). Adding (20) from $t=1$ to $T$ yields
$\text{Regret}_{T}$ $\leq\sum_{t=1}^{T}\frac{\eta_{t}}{2}\left\|\nabla f_{t}(x_{t})\right\|^{2}$
$\leq\frac{L^{2}}{2}\sum_{t=1}^{T}\frac{1}{t}$
$=O(\log T),$
where the second inequality is due to Assumption 2 and (18).
It remains to account for $\text{CCV}_{T,2}$. For this, concatenate Lemma 3 with (18) to get $\text{CCV}_{T,2}\leq O(\log T)$, thereby closing the proof.
6 Extensions
We now briefly mention two straightforward extensions of CLASP: one to accommodate multiple constraints; the other to handle persistent constraints.
Multiple dynamic constraints.
In deriving CLASP, Algorithm 1, we assumed that only one constraint function $g_{t}$ is revealed at iteration $t$. In some applications, however, several constraints might be revealed, say, $M$ constraint functions, denoted $g_{t;m}$ for $1\leq m\leq M$. Accordingly, CCV might be adjusted to
$\text{CCV}^{1}_{T,2}=\sum_{t=1}^{T}\sum_{m=1}^{M}\left(g_{t;m}^{+}(x_{t})\right)^{2}\quad\text{or}\quad\text{CCV}^{\infty}_{T,2}=\sum_{t=1}^{T}\max\left\{(g_{t;m}^{+}(x_{t}))^{2}\,\colon\,1\leq m\leq M\right\}.$
Adapting CLASP to this setting is easy: it suffices to define the set ${\mathcal{K}}_{t}$ as ${\mathcal{K}}\cap C_{t;m_{t}}$, where $C_{t;m}=\left\{x\in\mathbb{R}^{n}\,\colon\,g_{t;m}(x_{t})\leq 0\right\}$ and $m_{t}$ denotes the index of the constraint incurring the largest violation at $x_{t}$, and correspondingly upgrade Assumption 3 so as to suppose the existence of $x_{T}^{\star}\in{\mathcal{K}}\cap{\mathcal{C}}_{T}$, with ${\mathcal{C}}_{T}=\cap_{t=1}^{T}\cap_{m=1}^{M}C_{t;m}$. It can be readily verified that the proofs of Lemmas 1 to 3 remain valid. In particular, Lemma 3 now asserts that $\sum_{t=1}^{T}(g_{t;m}^{+}(x_{t}))^{2}\leq O\left(\sum_{t=1}^{T}\eta_{t}\right)$ for any fixed $m\in\{1,\ldots,M\}$. This and the fact that there is a finite number $M$ of constraints imply that the statement of Lemma 3 continues to hold when $\text{CCV}_{T,2}$ is replaced by $\text{CCV}^{1}_{T,2}$ or $\text{CCV}^{\infty}_{T,2}$. Owing to the modular structure of the CLASP proof, which treats regret and CCV separately, Theorems 1 and 2 remain valid as well. To conclude, the theoretical guarantees of CLASP are preserved.
Persistent constraints.
The canonical interpretation in the COCO literature about dynamic constraints is that they are transient: the constraint $g_{t}$ should be satisfied by decision $x_{t}$, but not necessarily by future decisions $x_{t+1},x_{t+2},\ldots$. This interpretation can be read in the common definitions of CCVs in that they penalize only the violation of $x_{t}$ induced by $g_{t}$—if that was not the case, if constraints were to be interpreted as persistent, then the definition of a CCV would have to keep the score of violations of $x_{t}$ induced by the history of constraints $g_{1},g_{2},\ldots,g_{t}$ so far, say, through the much more stricter metric
$\text{CCV}_{T,2}^{\text{hist}}=\sum_{t=1}^{T}\sum_{\tau=1}^{t}(g_{\tau}^{+}(x_{t}))^{2}.$ (21)
The underlying interpretation of dynamic constraints as being transient can also be inferred from standard COCO algorithms, in the sense that they generate the decision $x_{t}$ without worrying, let alone enforcing, that it satisfies previously revealed constraints (in this regard, the only exception that we know of is Vaze and Sinha (2025), though the authors do not articulate that such property is built-in in their algorithm to address explicitly persistent constraints).
In any case, although we are not aware of specific applications that are best modeled by persistent constraints, CLASP can be effortlessly adjusted to such settings if needed by defining ${\mathcal{K}}_{t}={\mathcal{K}}\cap\bigcap_{t=1}^{T}C_{t}$, with $C_{t}=\left\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\right\}$. This forces $x_{t}$ to comply with all constraints revealed so far, that is, $g_{\tau}(x_{t})\leq 0$ holds for $\tau<t$, in which case $\text{CCV}_{T,2}^{\text{hist}}$ collapses back to $\text{CCV}_{T,2}$, and the theoretical analysis goes through unchanged, keeping the guarantees of CLASP intact.
7 Numerical Experiments
In this section, we present a synthetic experiment to evaluate the performance of our proposed algorithm. We compare CLASP with the AdaGrad-based algorithm of Sinha and Vaze (2024), the RECOO algorithm of Guo et al. (2022), and the Switch algorithm of Vaze and Sinha (2025). We report results for the cumulative loss and for the violation metrics $\mathrm{CCV}_{T,1}$ and $\mathrm{CCV}_{T,2}$.
On evaluating CCVT,2.
In line with prior COCO work, we measure $\mathrm{CCV}_{T,2}$ a posteriori to illustrate how this theoretically motivated quantity behaves in practice.
Further synthetic and real-data experiments can be found in Appendix A.4. All results were averaged over 100 trials and are reported with 95% confidence intervals. The experiments were performed on a 2020 MacBook Air 13”, with an 8-core Apple M1 processor and 16 GB of RAM.
7.1 Online Linear Regression
This experiment is similar to the one presented by Guo et al. (2022) for the setting of adversarial constraints with a synthetic dataset. At each iteration $t$, the loss function is given by $f_{t}(x)=\|H_{t}\,x-y_{t}\|$, with $H_{t}\in\mathbb{R}^{4\times 10}$, $x\in\mathbb{R}^{10}$ and $y_{t}\in\mathbb{R}^{4}$, where $H_{t}(i,j)\sim U(-1,1)$, $1\leq i\leq 4$, $1\leq j\leq 10$, and $y_{t}(i)=H_{t}(i)\mathbf{1}+\epsilon_{i}$, such that $\mathbf{1}\coloneqq(1,\dots,1)\in\mathbb{R}^{10}$ and $\epsilon_{i}$ denotes the standard normal random variable. On the other hand, the constraint function is given by $g_{t}(x)=A_{t}\,x-b_{t}$, with $A_{t}\in\mathbb{R}^{4\times 10}$ and $b_{t}\in\mathbb{R}^{4}$, where $A_{t}(i,j)\sim U(0,2)$ and $b_{t}(i)\sim U(0,1)$, for $1\leq i\leq 4$ and $1\leq j\leq 10$. Note that $g_{t}$ is a vector-valued map taking values in $\mathbb{R}^{4}$, and thus corresponds to multiple constraints of the form $g_{t,m}(x)\leq 0$, for $1\leq m\leq 4$. However, we can replace it with a single scalar function, which is the pointwise maximum of the constraints $g_{t}(x)=\max\{g_{t,m}(x)\,\colon 1\leq m\leq 4\}$ (see Section 6 for details). The static decision set is given by $\mathcal{K}=\left\{x\in\mathbb{R}^{10}:0\leq x_{i}\leq 1,\mbox{for }i=1,\dots,10\right\}$. The vector $x_{1}\in\mathbb{R}^{n}$ in all algorithms was initialized such that $x_{i}\sim U(0,1)$, for $1\leq i\leq n$, where $U(a,b)$ stands for the uniform distribution supported on the interval $(a,b)$.
(a) Cumulative loss
(b) $\text{CCV}_{t,1}$
(c) $\text{CCV}_{t,2}$
Figure 1: Online linear regression with adversarial constraints. We report (a) cumulative loss, (b) linear violation $\mathrm{CCV}_{T,1}$, and (c) squared violation $\mathrm{CCV}_{T,2}$. AdaGrad achieves the lowest loss but at the cost of very large constraint violations. CLASP controls both violation metrics competitively with RECOO, while remaining more memory-efficient than Switch. Switch attains the smallest violations overall but with higher per-iteration complexity. All $\mathrm{CCV}_{T,2}$ values are reported a posteriori.
Figure 1 reports cumulative loss, linear violation ($\mathrm{CCV}_{T,1}$), and squared violation ($\mathrm{CCV}_{T,2}$) over iterations. The AdaGrad-based method of Sinha and Vaze (2024) attains the lowest cumulative loss, but only by incurring much larger constraint violations. By contrast, CLASP (orange) controls both $\mathrm{CCV}_{T,1}$ and $\mathrm{CCV}_{T,2}$ at levels comparable to RECOO (green), while remaining simpler and more memory-efficient. Switch (red) achieves the smallest overall violations, but at the cost of higher per-iteration complexity. Notably, on the squared violation metric $\mathrm{CCV}_{T,2}$—the focus of our analysis—CLASP is competitive with the best-performing baselines, confirming in practice that it can keep violation severity sublinear while maintaining reasonable regret.
8 Conclusions
We introduced CLASP, an online COCO algorithm that handles convex losses and dynamic constraints. CLASP aims at minimizing both loss regret and cumulative constraint violation (CCV). In this work, the metric is the squared Cumulative Constraint Violation, to account for large violations. For general convex losses, CLASP offers a tunable trade-off between regret and CCV, matching the best performance of previous works designed for static constraints. More importantly, for strongly convex losses, CLASP universally achieves logarithmic bounds on both regret and CCV—an advance that, to the extent of our knowledge, is established here for the first time. Algorithmically, CLASP consists of just a gradient step followed by a projection at each iteration, while its analysis is simplified by leveraging the firmly non-expansiveness property of projections. This yields a modular proof structure that disentangles the treatment of regret from that of CCV. This modularity, in turn, makes extensions of CLASP to other settings (e.g., multiple or persistent constraints) straightforward.
Appendix A Appendix
A.1 Proof of Lemma 1
Note that $x_{T}^{\star}\in\mathcal{K}\cap{\mathcal{C}}_{T}\subset{\mathcal{K}}_{t}$, implying $x_{T}^{\star}=P_{{\mathcal{K}}_{t}}(x_{T}^{\star})$ for $1\leq t\leq T$. We have
$\|x_{t+1}-x_{T}^{\star}\|^{2}$ $=\left\|P_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})-P_{{\mathcal{K}}_{t}}(x_{T}^{\star})\right\|^{2}$
$\leq\|\widehat{x}_{t+1}-x_{T}^{\star}\|^{2}-d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$
$=\|x_{t}-\eta_{t}\nabla f_{t}(x_{t})-x_{T}^{\star}\|^{2}-d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$
$=\|x_{t}-x_{T}^{\star}\|^{2}-2\eta_{t}\langle\nabla f_{t}(x_{t}),x_{t}-x_{T}^{\star}\rangle+\eta_{t}^{2}\|\nabla f_{t}(x_{t})\|^{2}-d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$
$\leq\|x_{t}-x_{T}^{\star}\|^{2}+2\eta_{t}\|\nabla f_{t}(x_{t})\|\|x_{t}-x_{T}^{\star}\|+\eta_{t}^{2}\|\nabla f_{t}(x_{t})\|^{2}-d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$
$\leq\|x_{t}-x_{T}^{\star}\|^{2}+2LD\eta_{t}+L^{2}\eta_{t}^{2}-d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2},$
where the first inequality is due to the firm non-expansiveness of the operator $P_{{\mathcal{K}}_{t}}$ (see (5)); the second inequality is the Cauchy-Schwarz inequality; and the third inequality follows from Assumptions 1 and 2.
Before proceeding, we would like to indicate the pivotal role played by the firmly non-expansiveness property of the operator $P_{{\mathcal{K}}_{t}}$. Indeed, if we used only its weaker non-expansiveness property, the right-hand side of the first inequality would read as $\left\|\widehat{x}_{t+1}-x_{T}^{\star}\right\|^{2}$, the key term $d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$ no longer present. The disappearance of this term would instantly sever the reasoning of the present lemma; as the subsequent lemmas depend upon it, the entire proof would be invalidated.
Re-arranging the last inequality yields
$d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}$ $\leq\|x_{t}-x_{T}^{\star}\|^{2}-\|x_{t+1}-x_{T}^{\star}\|^{2}+2LD\eta_{t}+L^{2}\eta_{t}^{2}$
$\leq\|x_{t}-x_{T}^{\star}\|^{2}-\|x_{t+1}-x_{T}^{\star}\|^{2}+(2LD+\theta L^{2})\eta_{t},$
where the last inequality is due to Assumption 4.
Summing up from $t=1$ to $T$, we obtain
$\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}\leq\|x_{1}-x_{T}^{\star}\|^{2}+\sum_{t=1}^{T}(2LD+\theta L^{2})\eta_{t}=O\left(\sum_{t=1}^{T}\eta_{t}\right).$
A.2 Proof of Lemma 2
We have
$d_{{\mathcal{K}}_{t}}(x_{t})$ $=d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1}+(x_{t}-\widehat{x}_{t+1}))$
$\leq d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})+\|x_{t}-\widehat{x}_{t+1}\|$
$=d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})+\eta_{t}\|\nabla f_{t}(x_{t})\|$
$\leq d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})+\eta_{t}L,$
where the first inequality follows from the Lipschitz continuity of the distance function $d_{{\mathcal{K}}_{t}}$ (see Section 2), and the second inequality from Assumption 2.
It follows that
$d_{{\mathcal{K}}_{t}}(x_{t})^{2}$ $\leq 2d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}+2\eta_{t}^{2}L^{2}$
$\leq 2d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}+2\eta_{t}\theta L^{2},$
where the first inequality comes from the general fact $(a+b)^{2}\leq 2a^{2}+2b^{2}$ for $a,b\in\mathbb{R}$, and the second inequality from Assumption 4.
Summing up from $t=1$ to $T$ and using Lemma 1, we obtain
$\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(x_{t})^{2}$ $\leq 2\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(\widehat{x}_{t+1})^{2}+2L^{2}\theta\sum_{t=1}^{T}\eta_{t}$
$\leq O\left(\sum_{t=1}^{T}\eta_{t}\right).$
A.3 Proof of Lemma 3
Note that $|g_{t}^{+}(u)-g_{t}^{+}(v)|\leq|g_{t}(u)-g_{t}(v)|$ for all $u,v\in\mathbb{R}^{n}$. Hence, $|g_{t}^{+}(u)-g_{t}^{+}(v)|\leq L\left\|u-v\right\|$ for all $u,v\in{\mathcal{K}}$ due to Assumption 3 and its consequence (9).
It follows that
$g_{t}^{+}(x_{t})$ $\leq g_{t}^{+}\left(P_{{\mathcal{K}}_{t}}(x_{t})\right)+L\left\|x_{t}-P_{{\mathcal{K}}_{t}}(x_{t})\right\|$
$=Ld_{{\mathcal{K}}_{t}}(x_{t}),$
since $g_{t}^{+}\left(P_{{\mathcal{K}}_{t}}(x_{t})\right)=0$ because $P_{{\mathcal{K}}_{t}}(x_{t})\in{\mathcal{K}}_{t}={\mathcal{K}}\cap\{x\in\mathbb{R}^{n}\,\colon\,g_{t}(x)\leq 0\}$ and any point $u\in{\mathcal{K}}_{t}$ satisfies $g_{t}(u)\leq 0$.
Equivalently, we have $g_{t}^{+}(x_{t})^{2}\leq L^{2}d_{{\mathcal{K}}_{t}}(x_{t})^{2}$. Summing up from $t=1$ to $T$ and using Lemma 2, we obtain
$\text{CCV}_{T,2}$ $=\sum_{t=1}^{T}(g_{t}^{+}(x_{t}))^{2}$
$\leq L^{2}\sum_{t=1}^{T}d_{{\mathcal{K}}_{t}}(x_{t})^{2}$
$\leq O\left(\sum_{t=1}^{T}\eta_{t}\right).$
A.4 Additional Numerical Experiments
In this section, we present additional numerical experiments. In particular, a version of the Online Linear Regression experiment with additional rounds, and an experiment for Online Support Vector Machine. The results were obtained by averaging over 100100 trials and reported with a $95\%$ confidence interval.
A.4.1 Online Linear Regression with additional rounds
We now present a version of the Online Linear Regression experiment with a larger number of rounds. As stated in Section 1.2, the Switch algorithm is memory-intensive as the set to which the algorithm performs the projections incorporates all previously revealed constraint functions. Therefore, as the iterations unfold, the set becomes increasingly more complex and the associated projection increasingly more expensive. So, we removed Switch from this experiment, as to be able to investigate a larger number of rounds.
(a) Cumulative loss
(b) $\text{CCV}_{t,1}$
(c) $CCV_{t,2}$
Figure 2: Online Linear Regression with additional rounds.
The results are similar to the ones in Section 7. However, while the Cumulative Constraint Violation (both $\text{CCV}_{T,1}$ and $\text{CCV}_{T,2}$) attained by CLASP and RECOO are very similar, note that, for the $\text{CCV}_{T,2}$ metric, for a larger number of rounds, the CLASP algorithm displays diminished variability.
A.4.2 Online Support Vector Machine
In our second experiment, we compare the performance of the algorithms for the online update of a Support Vector Machine (SVM) (Boser et al., 1992; Bishop and Nasrabadi, 2006). We consider that, at each iteration $t$, we receive a new labeled sample $(u_{t},v_{t})$, with $u_{t}\in\mathbb{R}^{P}$ the feature vector and $v_{t}\in\{\pm 1\}$ the label. Thus, at each round $t$, we can formulate the optimization problem in the COCO setting as
$\underset{x\coloneqq(w,b)\in\mathcal{K}}{\mbox{minimize}}\quad\frac{1}{2}\|w\|^{2},\quad\mbox{subject to}\quad-v_{t}\left(w^{T}u_{t}-b\right)+1\leq 0,$
or, in terms of the COCO framework, the revealed loss function is $f_{t}(x)=\frac{1}{2}\|w\|^{2}$ and the revealed constraint function is $g_{t}(x)=-v_{t}\left(w^{T}u_{t}-b\right)+1$. For this experiment, we use the real-world dataset about wine quality from their physicochemical properties (for the details about the dataset, see (Cortez et al., 2009)) [^1] This dataset is released under a Creative Commons Attribution 4.0 International ($C$ BY 4.0) license.. The dataset contains 6497 samples, each sample contains $P=11$ features, and the quality of the wine is classified between 1 and 9. Some of the features were in $g/dm^{3}$, while others were in $mg/dm^{3}$. We ensured that all density features were expressed in $g/dm^{3}$. In this experiment, we consider the binary classification setting, where we label with 11 the wine samples with quality equal to or greater than 7, and label with 1-1 the remaining. Thus, we want our classifier to distinguish between high-quality and low-quality wines. In each trial, we reshuffle the dataset so that the order of samples in each trial is always different. The vector $x_{1}\in\mathbb{R}^{P+1}$ for all algorithms was initialized such that $x_{i}\sim U(-1,1)$, for $1\leq i\leq P+1$. The results were obtained by averaging over 100100 trials and reported with a $95\%$ confidence interval.
From domain knowledge, we can bound the norm of the feature vector as each physicochemical property has an interval of possible values (for simplicity, we analyzed the possible values encountered in the dataset and concluded that $0\leq u_{t}\leq 70$ for all $t$, thus we can use this to define the Lipschitz constant of the constraint functions $L=70\sqrt{P}$). While the SVM is an unconstrained problem, our algorithms assume that the decision set is compact; thus, based on the constraints on the feature vectors, we define $\mathcal{K}=\{x=(w,b)\in\mathbb{R}^{P+1}:\|x\|\leq 70\sqrt{P}\}$.
(a) Cumulative loss
(b) $\text{CCV}_{t,1}$
(c) $\text{CCV}_{t,2}$
Figure 3: Online Support Vector Machine
Remark.
In this experiment, the feasibility assumption is not satisfied, i.e., the set $\mathcal{K}\cap\mathcal{C}_{T}=\varnothing$, where $\mathcal{C}_{T}=\bigcap_{t=1}^{T}C_{t}$. Although this property is used in the analysis of the algorithms in COCO to obtain regret and cumulative constraint violation bounds, most of the algorithms in COCO can still be applied to problems without this property. However, the Switch algorithm, due to the exploitation of nested convex bodies, cannot be applied when the feasibility property is not satisfied. Since, at each iteration, the algorithm projects onto the intersection of past constraint functions, then there will be some $T_{0}$, with $1\leq T_{0}\leq T$, such that $\mathcal{C}_{T_{0}}=\bigcap_{t=1}^{T_{0}}C_{t}=\varnothing$. Therefore, in this experiment, we cannot compare the performance of the Switch algorithm, but compare our algorithm CLASP with the AdaGrad and RECOO algorithms.
In Fig. 3, we can visualize both the cumulative loss and the CCV of the different algorithms for each round. In this experiment, the more informative results are in Figs. 3(b) and 3(c), as a constraint violation translates as a transgression of the margin. We see for both constraint violation metrics, $\text{CCV}_{T,1}$ and $\text{CCV}_{T,2}$, the CLASP algorithm achieves better results and with less variability. Furthermore, we see that for the metric $\text{CCV}_{T,2}$, the difference in performance is significantly better.